OpenStudy (goformit100):

If a bus moves 60 kmph it will have 3 hours delay. if it moves 90 kmph it will arrive 1 hour earlier. how many kilometers is this route?

OpenStudy (shubhamsrg):

distance/speed = time x/60 = t+3 x/90 = t-1 solve for x. ^_^

OpenStudy (goformit100):

Thankyou Sir

OpenStudy (shubhamsrg):

thank you @DLS

OpenStudy (dls):

@shubhamsrg is right,i inboxed him the answer,copy paste is legit.

OpenStudy (shubhamsrg):

I already said thank you! >.<

OpenStudy (dls):

:P